OpenStudy (hovlaus):

u u+1+ ----- u+1 to: (u+1)(u+1)+u ----------- u+1 How?

OpenStudy (anonymous):

\[ u+1 = \frac{u+1}{u+1}\cdot (u+1) = \frac{(u+1)^2}{u+1} \]

OpenStudy (anonymous):

So \[ u+1+\frac{u}{u+1} = \frac{(u+1)^2}{u+1}+\frac{u}{u+1} = \frac{(u+1)^2+u}{u+1} \]

OpenStudy (hovlaus):

Thank you very much, wio!